how many ml of soln are required to prepare 0.02 mol of 2.5m Nacl soln?


1
Expert's answer
2022-05-27T14:39:01-0400

CM — concentration of molarity (mol/l) = 2,5 mol/l;

n — amount of substance (mol) = 0,02 mol

V — volume (litr).


V = n / CM = 0,02 / 2,5 = 0,008 litr = 8 ml

ANSWER: 8 ml